OpenStudy (anonymous):

Solve and explain please ^u^ Factoring trinomials: 2s(squared) + 5s + 3 = ?

OpenStudy (turingtest):

factoring by grouping means we split up the middle term in a certain way it is hard to explain and easier to see if you watch carefully\[2s^2+5s+3=2s^2+2s+3s+3\]notice the way I split up the middle term we can now factor a 2s out of the first two terms, and a 3 on the next two:\[2s^2+2s+3s+3=2s(s+1)+3(s+1)\]and alost by magic now we can factor out s+1 from the above and get\[2s(s+1)+3(s+1)=(2s+3)(s+1)\]and that is factoring by grouping

OpenStudy (anonymous):

2s^2 + 5s + 3 (s + 1) (2s + 3) = 0 => s = -1, s = -3/2

OpenStudy (turingtest):

though we are not told that it equals zero...

OpenStudy (anonymous):

I thought the question is solve, so I assume solve for x :)